### v2

# Hoaxpedia: A Unified Wikipedia Hoax Articles Dataset

Abstract
Hoaxes are a recognised form of disinformation created deliberately, with potential serious implications in the credibility of reference knowledge resources such as Wikipedia. What makes detecting Wikipedia hoaxes hard is that they often are written according to the official style guidelines. In this work, we first provide a systematic analysis of the similarities and discrepancies between legitimate and hoax Wikipedia articles, and introduce Hoaxpedia, a collection of 311 Hoax articles (from existing literature as well as official Wikipedia lists) alongside semantically similar real articles. We report results of binary classification experiments in the task of predicting whether a Wikipedia article is real or hoax, and analyze several settings as well as a range of language models. Our results suggest that detecting deceitful content in Wikipedia based on content alone, despite not having been explored much in the past, is a promising direction. [^1] The Dataset(view-only, access upon request) is available in: https://osf.io/rce8m/?view_only=ed469941644c496fb4a6425297ced1f2. We will publicly release our models and the datasets in Huggingface upon acceptance
Hoaxpedia: A Unified Wikipedia Hoax Articles Dataset
Hsuvas Borkakoty1, Luis Espinosa-Anke1,2, 1Cardiff NLP, School of Computer Science and Informatics, Cardiff University, UK 2AMPLYFI, UK {borkakotyh,espinosaankel}@cardiff.ac.uk
1 Introduction
Wikipedia is, as Hovy et al. (2013) define it, the “largest and most popular collaborative and multilingual resource of world and linguistic knowledge”, and it is acknowledged that its accuracy is on par with or superior than, e.g., the Encyclopedia Britannica Giles (2005). However, as with any other platform, Wikipedia is also the target of online vandalism, and hoaxes, a more obscure, less obvious form of vandalism [^2] https://en.wikipedia.org/wiki/Wikipedia:Do_not_create_hoaxes., constitute a threat to the overall integrity of this collaborative encyclopedia (Kumar et al., 2016; Wong et al., 2021; Wang and McKeown, 2010), precisely because of its “publish first, ask questions later” policy Asthana and Halfaker (2018). Although Wikipedia employs community based New Page Patrol systems to check the credibility of a newly created article, the process is always in backlog [^3] https://en.wikipedia.org/wiki/Wikipedia:New_pages_patrol., making the process overwhelming Schneider et al. (2014).
Hoax articles, created to deliberately spread false information Kumar et al. (2016), harm the credibility of Wikipedia as a knowledge resource, and generate concerns among its users Hu et al. (2007). Since manual inspection of quality is a process typically in backlog Dang and Ignat (2016), the automatic detection of such articles is a desirable feature. However, most works in the literature have centered their efforts in metadata associated with hoax articles, e.g., user activity, appearance features or revision history Zeng et al. (2006); Elebiary and Ciampaglia (2023); Kumar et al. (2016); Wong et al. (2021); Hu et al. (2007); Susuri et al. (2017). For example, Adler et al. (2011) introduced a vandalism detection system using metadata, content and author reputation features, whereas Kumar et al. (2016) provides a comprehensive study of hoax articles and their timeline from discovery to deletion. In this work, the authors define the characteristics of a successful hoax, with a data-driven approach based on studying a dataset of 64 articles (both hoax and real), on top of which they train statistical classifiers. Furthermore, other works have compared network traffic and features of hoax articles to those of other articles published the same day (Elebiary and Ciampaglia, 2023), and conclude that hoax articles attract more attention after creation than cohort articles. Finally, Wong et al. (2021) study various Wikipedia vandalism types and introduce the Wiki-Reliability dataset, which comprises articles based on 41 author-compiled templates. This dataset contains 1,300 articles marked as hoax, which are real articles with false information, a.k.a hoax facts (Kumar et al., 2016).
We part ways from previous works and focus exclusively on the content of hoax articles, and aim to answer the research question “Can hoax articles be distinguished from real articles using NLP techniques by looking exclusively at the article’s content”? We first construct a dataset (HoaxPedia) containing 311 hoax articles and around 30,000 plausible negative examples, i.e., real Wikipedia articles that are semantically similar to hoax articles, to create a set of negatives that cover similar topics to hoax articles (e.g., a newly discovered species). We also explore whether a Wikipedia definition (the first sentence of the article) can provide hints towards its veracity. Our results (reported at different ratios of hoax vs real articles) suggest that, while style and shallow features are certainly not good predictors of hoax vs real Wikipedia articles, LMs are capable of exploiting other more intricate features, and open a promising research direction focused on content-based hoax flagging. Our contributions in this work can be summarised as follows.
  • We systematically contrast a set of proven Wikipedia hoax articles vs. legitimate articles.
  • We propose HoaxPedia, a novel Wikipedia Hoax article dataset with 311 hoax articles and semantically similar legitimate articles collected from Wikipedia.
  • We conduct binary classification experiments on this dataset, using a range of language models to accurately predict whether an article is a hoax or real based solely on its content.
2 HoaxPedia Construction
HoaxPedia is constructed by unifying five different resources that contain known hoaxes, e.g., from Kumar et al. (2016); Elebiary and Ciampaglia (2023), as well as the official Wikipedia hoaxes list [^4] https://en.wikipedia.org/wiki/Wikipedia:List_of_hoaxes_on_Wikipedia and the Internet Archive. We used Internet Archive to manually retrieve Wikipedia pages that are now deleted from Wikipedia, but were at one point in the past identified as hoaxes. We manually verify each of the article we collect from Wikipedia and Internet Archive as hoax using their accompanied deletion discussion and reasons for citing them as a hoax. In terms of negative examples, while we could have randomly sampled Wikipedia pages, this could have introduced a number of biases in the dataset, e.g., hoax articles contain historical events, personalities or artifacts, and thus we are interested in capturing a similar breadth of topics, entities and sectors in the negative examples so that a classifier cannot rely on these spurious features. We select these negative examples (truthful articles) by ensuring they correspond to authentic content. This is achieved by verifying they do not carry the Db-hoax flag, which Wikipedia’s New Page Patrol policy uses to mark potential hoaxes. Within this set, we extract negative examples as follows. Let $H$ be the set of hoax articles, and $W$ the set of candidate real Wikipedia pages, with $T_{H}=\{t_{H^{1}},\dots,t_{H^{p}}\}$ and $T_{W}=\{t_{W^{1}},\dots,t_{W^{q}}\}$ their corresponding vector representations, and $p$ and $q$ the number of hoax and candidate Wikipedia articles, respectively. Then, for each SBERT(all-MiniLM-L6-v2) ( (Reimers and Gurevych, 2019) title embedding $t_{H^{i}}\in T_{H}$, we retrieve its top $k$ nearest neighbors (NN) from $T_{W}$ via cosine similarity cos. We experiment with different values for $k$, specifically $k\in\{2,10,100\}$:
$\mbox{{NN}}\left(t_{H^{i}}\right)=\{t_{W^{j}}:j\in J_{k}(t_{H^{i}})\}$
where $J_{k}(t_{H^{i}})$ contains the top $k$ cosine similarities in $T_{W}$ for a given $t_{H^{i}}$, and
$\mbox{{cos}}\left(t_{H^{i}},t_{W^{j}}\right)=\frac{t_{H^{i}}\cdot t_{W^{j}}}{||t_{H^{i}}||||t_{W^{j}}||}$
The result of this process is a set of positive (hoax) articles and a set of negative examples we argue will be similar in content and topic, effectively removing any topic bias from the dataset.
3 Hoax vs. Real, a Surface-Level Comparison
To maintain the longevity and to avoid detection, hoax articles follow Wikipedia guidelines and article structure. This raises the following question: “how (dis)similar are hoaxes with respect to a hypothetical real counterpart?”. Upon inspection, we found comments in the deletion discussions such as “I wouldn’t have questioned it had I come across it organically” (for the hoax article The Heat is On  [^5] https://en.wikipedia.org/wiki/Wikipedia:Articles_for_deletion/The_Heat_Is_On_(TV_series)), or “The story may have a "credible feel” to it, but it lacks any sources”, a comment on article Chu Chi Zui [^6] https://en.wikipedia.org/wiki/Wikipedia:Articles_for_deletion/Chu_Chi_Zui. Comments like these highlight that hoaxes are generally well written (following Wikipedia’s guidelines), and so we proceed to quantify their stylistic differences in a comparative analysis that looks at: (1) article text length comparison; (2) sentence and word length comparison; and (3) a readability analysis.
Article Text length distribution:
Following the works of Kumar et al. (2016), we conduct a text length distribution analysis with hoax and real articles, and verify they show a similar pattern (as shown in Figure 1), with similar medians for hoax and real articles, specifically 1,057 and 1,777 words, respectively.
Figure 1: Text length distribution for hoax and real articles.
Average sentence and word length:
Calculating average sentence and word length for hoaxes and real articles separately can be a valuable proxy for identifying any obvious stylistic or linguistic (e.g., syntactic complexity) patterns. We visualize these in a series of box plots in Figure 2. They clearly show a similar style, with sentence and word length medians at 21.23 and 22.0, and 4.36 and 4.35 for real and hoax articles respectively.
(a) Flesch-Kincaid Grade
(b) Average Sentence Length
(c) Average Word Length
Figure 2: Results of different stylistic analyses on Hoax (red) and Real (blue) articles.
Readability Analysis:
Readability analysis gives a quantifiable measure of the complexities of texts. It can clarify on easement of understanding the text, revealing patterns that can be used to either disguise disinformation like hoaxes or convey clear, factual content. For readability analysis, we use Flesch-Kincaid (FK) Grading system (Flesch, 2007), a metric that indicates the comprehension difficulty when reading a passage in the context of contemporary academic English. This metric gives us an aggregate of the complexity of documents, their sentences and words, measured by the average number of words per sentence and the average number of syllables per word. After obtaining an average for both hoax and real articles, we visualize these averages again in Figure 2, we find a median of 9.4 for real articles and 9.5 for hoax articles, again highlighting the similarities between these articles.
The above analysis suggests that hoax Wikipedia articles are indeed well disguised, and so in the next section we propose a suite of experiments for hoax detection based on language models, setting an initial set of baselines for this novel dataset.
4 Experiments
We cast the problem of identifying hoax vs. legitimate articles as a binary classification problem, in whcih we evaluate a suite of LMs, specifically: BERT-family of models (BERT-base and large Devlin et al. (2019), RoBERTa-base and large Liu et al. (2019), Albert-base and large Lan et al. (2019)), as well as T5 (Base and Large) Raffel et al. (2020) and Longformer (Base) Beltagy et al. (2020). We use the same training configuration for the BERT-family of models, T5 models and Longformer, and set the generation objective as Binary classification for the T5 models. In terms of data size, we consider the three different scenarios outlined in Section 2 (2x, 10x and 100x negative examples). This approach naturally increases the challenge for the classifiers. The details about the data used in different settings are given in Appendix A.
In addition to the three different settings for positive vs negative ratios, we also explore how much text is actually needed to catch a hoax, or in other words, are definition sentences in hoax articles giving something away? This is explored by running our experiments on the full Wikipedia articles, on one hand, and on the definition (first sentence alone), on the other. This latter setting is interesting from a lexicographic perspective because it helps us understand if the Wikipedia definitions show any pattern that a model could exploit. Moreover, from the practical point of view of building a classifier that could dynamically “patrol” Wikipedia and flag content automatically, a definition-only model would be more interpretable (with reduced amiguity and focusing on core meaning/properties of the entity) and could have less parameters (handling smaller vocabularies, and compressed knowledge), which would have practical retraining/deployment implications in cost and turnaround.
5 Results
Our experiments are aimed to explore the impact of data imbalance and content length. In both cases, we compare several classifiers and analyze whether model size (in number of parameters) is correlated with performance. In terms of evaluation metrics, all results we report are F1 on the positive class (hoax). In definition-only setting, we find that models evaluated on datasets that are relatively balanced (2 Real articles for every hoax) show a stable performance, but they degrade drastically as the imbalance increases. In terms of robustness, RoBERTa is the most consistent, with an F1 of around 0.6 for all three settings, whereas Albert models perform poorly (exhibiting, however, some interesting behaviours, which we will discuss further). For the full text setting, we find that Longformer models performs well, with an F1 of 0.8. Surprisingly, the largest model we evaluated (T5-large) is not the best performing model, although this could point to under-fitting, the dataset potentially being too small for a model this size to train properly. Another interesting behaviour of T5-large is that in the 1 Hoax vs 2 Real setting, it shows exactly the same performance, whether seeing a definition or the full text. On the other side of the spectrum, we find that Albert models are the ones showing the highest improvement when going from definition to full text. This is interesting, as it shows a small model may miss nuances in definitions but can still compete with, or even outperform, larger models.
Definition Fulltext
Model Model Size 1H2R 1H10R 1H100R 1H2R 1H10R 1H100R
Albert-base-v2 12M 0.23 0.17 0.06 0.67 0.47 0.11
Albert-large-v2 18M 0.28 0.30 0.15 0.72 0.63 0.30
BERT-base 110M 0.42 0.30 0.14 0.55 0.57 0.32
RoBERTa Base 123M 0.57 0.59 0.53 0.82 0.75 0.63
Longformer-base 149M 0.43 0.35 0.54 0.80 0.78 0.67
T5-Base 220M 0.48 0.25 0.14 0.51 0.27 0.23
BERT-large 340M 0.43 0.36 0.17 0.61 0.64 0.33
RoBERTa-large 354M 0.58 0.63 0.62 0.84 0.81 0.79
T5-large 770M 0.54 0.32 0.13 0.54 0.43 0.37
Table 1: F1 on the positive class - hoax at different degrees of data imbalance for Definition and Fulltext Settings(H: Hoax, R: Real)
A perhaps not too surprising observation is that all models improve after being exposed to more text, as seen in Table 1, increasing their F1 about 20% on average, and sometimes even up to 30. This confirms that definitions alone are not a sufficiently strong signal for detecting hoax articles, although there are notable exceptions. Moreover, in terms of absolute performance, RoBERTa models perform decently, although significantly below their full text settings. It is interesting to note that Longformer base yields much better results in the hardest setting (1 Hoax vs 100 Real) when exposed only to definitions. This is indeed a surprising and counter intuitive result that deserves future investigation.
5.1 Ablation: Effect of Definitions on Classifying Hoaxes
We run a data ablation test on the full-text split of the dataset to find the impact of definition sentences. To this end, we remove the first sentence of each article, and replicate the “full text” classification experiment, focusing on RoBERTa-Large, the most consistent model. The results of this ablation experiment are shown in Table 2,suggesting that F1 goes down about 2% for the positive class when the definition sentence is missing. This shows that definitions show critical information about entities and events in Wikipedia, but often are not the place where hoax features would emerge, and therefore removing them from the full text doesn’t change much the story.
Model Setting Precision Recall F1
RoBERTaL Ft 1:2 0.83 0.80 0.82
RoBERTaL Ft 1:10 0.82 0.71 0.76
RoBERTaL Ft 1:100 0.67 0.51 0.58
Table 2: Performance of RoBERTa-Large on Binary Classifications Without Definition Sentences in Articles (with Fulltext Hoax:Real ratio in Settings column)
6 Conclusion and Future Work
We have introduced HoaxPedia, a dataset containing hoax articles extracted from Wikipedia, from a number of sources, from official lists of hoaxes, existing datasets and Web Archive. We paired these hoax articles with similar real articles, and after analyzing their main properties (concluding they are written with very similar style and content), we report the results of a number of binary classification experiments, where we explore the impact of (1) positive to negative ratio; and (2) going from the whole article to only the definition. This is different from previous works in that we have exclusively looked at the content of these hoax articles, rather than metadata such as traffic or longevity. For the future, we would like to further refine what are the criteria used by Wikipedia editors to detect hoax articles, and turn those insights into a ML model, and explore other types of non-obvious online vandalism.
7 Ethics Statement
This paper is in the area of online vandalism and disinformation detection, hence a sensitive topic. All data and code will be made publicly available to contribute to the advancement of the field. However, we acknowledge that deceitful content can be also used with malicious intents, and we will make it clear in any associated documentation that any dataset or model released as a result of this paper should be used for ensuring a more transparent and trustworthy Internet.
Appendix A Appendix: Dataset Details
We release our dataset in 3 settings as mentioned in Section 4. The settings with data splits and their corresponding sizes are mentioned in Table 3.
Number of Instances
Dataset Setting Dataset Type Split Non-hoax Hoax Total
1Hoax2Real Definition Train 426 206 632
Test 179 93 272
1Hoax2Real Full Text Train 456 232 688
Test 200 96 296
1Hoax10Real Definition Train 2225 203 2428
Test 940 104 1044
1Hoax10Real Full Text Train 2306 218 2524
Test 973 110 1083
1Hoax100Real Definition Train 20419 217 20636
Test 8761 82 8843
1Hoax100Real Full Text Train 22274 222 22496
Test 9534 106 9640
Table 3: Dataset details in different settings and splits
Appendix B Appendix: Training Details
We train our models with the configuration given below. We use one NVIDIA RTX4090, one NVIDIA V100 and one NVIDIA A100 GPU to train the models.
  • Learning rate: 2e-06
  • Batch size: 4 (for Fulltext experiments) and 8 (For Definition experiments)
  • Epochs: 30
  • Loss Function: Weighted Cross Entropy Loss
  • Gradient Accumulation Steps: 4
  • Warm-up steps: 100

### v3

@@
 What makes detecting Wikipedia hoaxes hard is that they often are written according to the official style guidelines.
-In this work, we first provide a systematic analysis of the similarities and discrepancies between legitimate and hoax Wikipedia articles, and introduce Hoaxpedia, a collection of 311 Hoax articles (from existing literature as well as official Wikipedia lists) alongside semantically similar real articles.
-We report results of binary classification experiments in the task of predicting whether a Wikipedia article is real or hoax, and analyze several settings as well as a range of language models.
-Our results suggest that detecting deceitful content in Wikipedia based on content alone, despite not having been explored much in the past, is a promising direction.
-[^1] The Dataset(view-only, access upon request) is available in: https://osf.io/rce8m/?view_only=ed469941644c496fb4a6425297ced1f2.
-We will publicly release our models and the datasets in Huggingface upon acceptance
+In this work, we first provide a systematic analysis of similarities and discrepancies between legitimate and hoax Wikipedia articles, and introduce Hoaxpedia, a collection of 311 hoax articles (from existing literature and official Wikipedia lists), together with semantically similar legitimate articles, which together form a binary text classification dataset aimed at fostering research in automated hoax detection.
+In this paper, We report results after analyzing several language models, hoax-to-legit ratios, and the amount of text classifiers are exposed to (full article vs the article’s definition alone).
+Our results suggest that detecting deceitful content in Wikipedia based on content alone is hard but feasible, and complement our analysis with a study on the differences in distributions in edit histories, and find that looking at this feature yields better classification results than context.
+[^1] The Dataset is available at: https://huggingface.co/datasets/hsuvaskakoty/hoaxpedia and associated codes are available at: https://github.com/hsuvas/hoaxpedia_dataset.git
 Hoaxpedia: A Unified Wikipedia Hoax Articles Dataset
-Hsuvas Borkakoty1, Luis Espinosa-Anke1,2, 1Cardiff NLP, School of Computer Science and Informatics, Cardiff University, UK 2AMPLYFI, UK {borkakotyh,espinosaankel}@cardiff.ac.uk
+Hsuvas Borkakoty1, Luis Espinosa-Anke1,2 1Cardiff NLP, School of Computer Science and Informatics, Cardiff University, UK 2AMPLYFI, UK {borkakotyh,espinosaankel}@cardiff.ac.uk
 1 Introduction
 Wikipedia is, as Hovy et al.
-(2013) define it, the “largest and most popular collaborative and multilingual resource of world and linguistic knowledge”, and it is acknowledged that its accuracy is on par with or superior than, e.g., the Encyclopedia Britannica Giles (2005).
-However, as with any other platform, Wikipedia is also the target of online vandalism, and hoaxes, a more obscure, less obvious form of vandalism [^2] https://en.wikipedia.org/wiki/Wikipedia:Do_not_create_hoaxes., constitute a threat to the overall integrity of this collaborative encyclopedia (Kumar et al., 2016; Wong et al., 2021; Wang and McKeown, 2010), precisely because of its “publish first, ask questions later” policy Asthana and Halfaker (2018).
-Although Wikipedia employs community based New Page Patrol systems to check the credibility of a newly created article, the process is always in backlog [^3] https://en.wikipedia.org/wiki/Wikipedia:New_pages_patrol., making the process overwhelming Schneider et al.
+(2013) defines it, the “largest and most popular collaborative and multilingual resource of world and linguistic knowledge”, and it is acknowledged that its accuracy is on par with or superior than, e.g., the Encyclopedia Britannica Giles (2005).
+However, as with any other platform, Wikipedia is also the target of online vandalism, and hoaxes, a more obscure, less obvious form of vandalism [^2] https://en.wikipedia.org/wiki/Wikipedia:Do_not_create_hoaxes., constitute a significant threat to Wikipedia’s overall integrity (Kumar et al., 2016; Wong et al., 2021; Wang and McKeown, 2010), among others, because of its “publish first, ask questions later” policy Asthana and Halfaker (2018).
+Although Wikipedia employs community based New Page Patrol systems to check the credibility of a newly created article, the process is always in backlog [^3] https://en.wikipedia.org/wiki/Wikipedia:New_pages_patrol., making it overwhelming Schneider et al.
 (2014).
-Hoax articles, created to deliberately spread false information Kumar et al.
-(2016), harm the credibility of Wikipedia as a knowledge resource, and generate concerns among its users Hu et al.
+Figure 1: Example of Hoax and Legitimate(Real) article
+Hoax articles (as shown in Figure 1), are created to deliberately spread false information Kumar et al.
+(2016), harm the credibility of Wikipedia as a knowledge resource and generate concerns among its users Hu et al.
 (2007).
@@
 (2011) introduced a vandalism detection system using metadata, content and author reputation features, whereas Kumar et al.
-(2016) provides a comprehensive study of hoax articles and their timeline from discovery to deletion.
-In this work, the authors define the characteristics of a successful hoax, with a data-driven approach based on studying a dataset of 64 articles (both hoax and real), on top of which they train statistical classifiers.
-Furthermore, other works have compared network traffic and features of hoax articles to those of other articles published the same day (Elebiary and Ciampaglia, 2023), and conclude that hoax articles attract more attention after creation than cohort articles.
+(2016) provide a comprehensive study of hoax articles and their timeline from discovery to deletion.
+In their work, the authors define the characteristics of a successful hoax, with a data-driven approach based on studying a dataset of 64 articles (both hoax and legitimate), on top of which they train statistical classifiers.
+Furthermore, other works have compared network traffic and features of hoax articles to those of other articles published the same day (Elebiary and Ciampaglia, 2023), and conclude that hoax articles attract more attention after creation than cohort (or legitimate) articles.
 Finally, Wong et al.
 (2021) study various Wikipedia vandalism types and introduce the Wiki-Reliability dataset, which comprises articles based on 41 author-compiled templates.
-This dataset contains 1,300 articles marked as hoax, which are real articles with false information, a.k.a hoax facts (Kumar et al., 2016).
-We part ways from previous works and focus exclusively on the content of hoax articles, and aim to answer the research question “Can hoax articles be distinguished from real articles using NLP techniques by looking exclusively at the article’s content”?
-We first construct a dataset (HoaxPedia) containing 311 hoax articles and around 30,000 plausible negative examples, i.e., real Wikipedia articles that are semantically similar to hoax articles, to create a set of negatives that cover similar topics to hoax articles (e.g., a newly discovered species).
+This dataset contains 1,300 articles marked as hoax, which are legitimate articles with false information, a.k.a hoax facts (Kumar et al., 2016).
+In this paper, we propose to incorporate content to the mix of features/signals that are fed to classifiers for the hoax detection task, and aim to answer the research question.
+We first construct a dataset (HoaxPedia) containing 311 hoax articles and around 30,000 plausible negative examples, i.e., legitimate Wikipedia articles that are semantically similar to hoax articles, to create a set of negatives that cover similar topics to hoax articles (e.g., a newly discovered species).
 We also explore whether a Wikipedia definition (the first sentence of the article) can provide hints towards its veracity.
-Our results (reported at different ratios of hoax vs real articles) suggest that, while style and shallow features are certainly not good predictors of hoax vs real Wikipedia articles, LMs are capable of exploiting other more intricate features, and open a promising research direction focused on content-based hoax flagging.
+Our results (reported at different ratios of hoax vs.
+legitimate articles) suggest that, while style and shallow features are certainly not good the best predictors, language models (LMs) are capable of exploiting other more intricate features (e.g., an article’s revision history) and open a promising research direction focused on content-based hoax flagging.
 Our contributions in this work can be summarised as follows.
-  • We systematically contrast a set of proven Wikipedia hoax articles vs. legitimate articles.
-  • We propose HoaxPedia, a novel Wikipedia Hoax article dataset with 311 hoax articles and semantically similar legitimate articles collected from Wikipedia.
-  • We conduct binary classification experiments on this dataset, using a range of language models to accurately predict whether an article is a hoax or real based solely on its content.
+  • We systematically contrast a set of proven Wikipedia hoax articles with legitimate articles.
+  • We propose HoaxPedia, a novel Wikipedia Hoax article dataset with 311 hoax articles and 30,000 semantically similar legitimate articles collected from Wikipedia.
+  • We conduct binary classification experiments on HoaxPedia, using a range of language models (including LLMs), features and hoax-to-legitimate ratio.
+2 Related work
+In what follows, we give a brief overview of disinformation detection, the datasets available for the community and the role of Wikipedia in disinformation detection, as our work falls in the intersection between disinformation detection and Wikipedia research.
+Disinformation detection and datasets:
+Disinformation and misinformation are two types of false information, they differ in that misinformation is inaccurate information created or propagated unknowingly, whereas disinformation is inaccurate information deliberately created to mislead the intended consumer Hernon (1995); Fallis (2014); Kumar et al.
+(2016); Ireton and Posetti (2018).
+Nonetheless, both are harmful towards information quality and reliability, thus posing risks toward different aspects of society Su et al.
+(2020).
+Alam et al.
+(2021) surveys disinformation detection from a multi-modal perspective, showing that it can come in different modalities like text, images, audio and video, with text being the most common.
+Datasets used for disinformation detection can be divided based on the length of input or claim: short sentences (such as tweets or Reddit posts) vs articles (common type being news articles), where most of the datasets follow claim-evidence based format Su et al.
+(2020).
+The short sentences or claim based datasets are mostly sourced from social media, such as X (formerly Twitter) Castillo et al.
+(2011); Derczynski et al.
+(2017); Zubiaga et al.
+(2018); López and Madhyastha (2021), Reddit Gorrell et al.
+(2018); Qu et al.
+(2022), or fact checking websites like Politifact [^4] https://www.politifact.com/ Wang (2017), Snopes [^5] https://www.snopes.com/ Vo and Lee (2020), or a combination of different fact checking websites Augenstein et al.
+(2019).
+These datasets contain claims which need to be verified as well as the response, in the form of True/False labeling backed by evidences, which are also provided in these datasets.
+Article level datasets, on the other hand, are varied, and focus on state-backed propaganda Heppell et al.
+(2023), German multi-label disinformation (the GerDISDETECT dataset) Schütz et al.
+(2024), or narratives at conflict dataset containing news articles Sinelnik and Hovy (2024), which mostly focuses on news article or propaganda based disinformation spreading.
+The datasets mentioned above are specialized towards topic/trend based or news based disinformation, lacking a generalised aspect of disinformation detection with free text.
+Wikipedia in disinformation detection:
+Wikipedia, as described by McDowell and Vetter (2020), serves as a source of information validation as backed by its large set of articles contributed by community.
+This is seen in action for fact verification task datasets such as FEVER Thorne et al.
+(2018b), TabFactA Chen et al.
+(2019), FNC-1 (Fake News Challenge-1) dataset Pomerleau and Rao (2017) etc.
+In these datasets, evidences for claims are collected from Wikipedia articles (eg.
+FEVER, FNC-1) and tables (eg.
+TabFactA).
+However, being a product of community effort, Wikipedia is also prone to vandalism and inaccurate contents McDowell and Vetter (2020), and the community outlines different policies to combat these issues [^6] https://en.wikipedia.org/wiki/Wikipedia:Vandalism.
+We also find efforts to automatize the process of detecting vandalism contents from Natural Language Processing perspective.
+Preliminary approaches include using feature based approaches extracted from metadata and editor behaviour to detect vandalism Wu et al.
+(2010); Javanmardi et al.
+(2011); Heindorf et al.
+(2016).
+Implementation of early warning system based on metadata and editor behavior is found in work of Kumar et al.
+(2015), where they propose a dataset of page metadata with list of blocked user and unblocked users, and proposes a feature based matrix for classification using auto-encoders.
+Yuan et al.
+(2017) proposes an edit history based approach, where they use behaviour of users over time as feature to create the embedding space for multi-source LSTM network Hochreiter and Schmidhuber (1997).
+However, these approaches do not consider article text as a marker to detect the vandalism.
+Wikipedia marks hoax articles as form of vandalism Kumar et al.
+(2016); Thorne et al.
+(2018a).
+A comprehensive discussion about different approaches for hoax detection are mentioned in Section 1.
+However, it is worth noting that existing research on this topic is not sufficient to make meaningful progress towards effective disinformation detection using Wikipedia, most of which can be attributed to unavailability of a dedicated dataset containing hoax articles as well as their metadata.
+We find a research gap towards developing datasets and tools for hoax detection where article contents are used, with foundation stone laid by works of Kumar et al.
+(2016) and Wong et al.
+(2021).
+Through our work, we aim to address this by introducing Hoaxpedia, where we unify the existing datasets we found based on article text, as well as collect articles from Wikipedia’s list of hoaxes.
 3 HoaxPedia Construction
@@
 In terms of negative examples, while we could have randomly sampled Wikipedia pages, this could have introduced a number of biases in the dataset, e.g., hoax articles contain historical events, personalities or artifacts, and thus we are interested in capturing a similar breadth of topics, entities and sectors in the negative examples so that a classifier cannot rely on these spurious features.
-We select these negative examples (truthful articles) by ensuring they correspond to authentic content.
+We select these negative examples by ensuring they correspond to authentic content.
 This is achieved by verifying they do not carry the Db-hoax flag, which Wikipedia’s New Page Patrol policy uses to mark potential hoaxes.
 Within this set, we extract negative examples as follows.
-Let $H$ be the set of hoax articles, and $W$ the set of candidate real Wikipedia pages, with $T_{H}=\{t_{H^{1}},\dots,t_{H^{p}}\}$ and $T_{W}=\{t_{W^{1}},\dots,t_{W^{q}}\}$ their corresponding vector representations, and $p$ and $q$ the number of hoax and candidate Wikipedia articles, respectively.
-Then, for each SBERT(all-MiniLM-L6-v2) ( (Reimers and Gurevych, 2019) title embedding $t_{H^{i}}\in T_{H}$, we retrieve its top $k$ nearest neighbors (NN) from $T_{W}$ via cosine similarity cos.
+Let $H$ be the set of hoax articles, and $W$ the set of candidate legitimate Wikipedia pages, with $T_{H}=\{t_{H^{1}},\dots,t_{H^{p}}\}$ and $T_{W}=\{t_{W^{1}},\dots,t_{W^{q}}\}$ their corresponding vector representations, and $p$ and $q$ the number of hoax and candidate Wikipedia articles, respectively.
+Then, for each SBERT (all-MiniLM-L6-v2) (Reimers and Gurevych, 2019) title embedding $t_{H^{i}}\in T_{H}$, we retrieve its top $k$ nearest neighbors (NN) from $T_{W}$ via cosine similarity cos.
 We experiment with different values for $k$, specifically $k\in\{2,10,100\}$:
@@
 The result of this process is a set of positive (hoax) articles and a set of negative examples we argue will be similar in content and topic, effectively removing any topic bias from the dataset.
-3 Hoax vs. Real, a Surface-Level Comparison
-To maintain the longevity and to avoid detection, hoax articles follow Wikipedia guidelines and article structure.
-This raises the following question: “how (dis)similar are hoaxes with respect to a hypothetical real counterpart?”.
+4 Hoax vs. legitimate, a Surface-Level Comparison
+To maintain longevity and avoid detection, hoax articles follow Wikipedia guidelines and article structure.
+This raises the following question: “how (dis)similar are hoaxes with respect to a hypothetical legitimate counterpart?”.
 Upon inspection, we found comments in the deletion discussions such as “I wouldn’t have questioned it had I come across it organically” (for the hoax article The Heat is On  [^8] https://en.wikipedia.org/wiki/Wikipedia:Articles_for_deletion/The_Heat_Is_On_(TV_series)), or “The story may have a "credible feel” to it, but it lacks any sources”, a comment on article Chu Chi Zui [^9] https://en.wikipedia.org/wiki/Wikipedia:Articles_for_deletion/Chu_Chi_Zui.
@@
 Following the works of Kumar et al.
-(2016), we conduct a text length distribution analysis with hoax and real articles, and verify they show a similar pattern (as shown in Figure 1), with similar medians for hoax and real articles, specifically 1,057 and 1,777 words, respectively.
-Figure 1: Text length distribution for hoax and real articles.
+(2016), we conduct a text length distribution analysis with hoax and legitimate articles, and verify they show a similar pattern (as shown in Figure 2), with similar medians for hoax and legitimate articles, specifically 1,057 and 1,777 words, respectively.
+Figure 2: Text length distribution for hoax and legitimate articles.
 Average sentence and word length:
-Calculating average sentence and word length for hoaxes and real articles separately can be a valuable proxy for identifying any obvious stylistic or linguistic (e.g., syntactic complexity) patterns.
+Calculating average sentence and word length for hoaxes and legitimate articles separately can be a valuable proxy for identifying any obvious stylistic or linguistic (e.g., syntactic complexity) patterns.
 We visualize these in a series of box plots in Figure 3.
-They clearly show a similar style, with sentence and word length medians at 21.23 and 22.0, and 4.36 and 4.35 for real and hoax articles respectively.
+They clearly show a similar style, with sentence and word length medians at 21.23 and 22.0, and 4.36 and 4.35 for legitimate and hoax articles respectively.
 (a) Flesch-Kincaid Grade
@@
 (c) Average Word Length
-Figure 2: Results of different stylistic analyses on Hoax (red) and Real (blue) articles.
+Figure 3: Results of different stylistic analyses on Hoax (red) and legitimate (blue) articles.
 Readability Analysis:
@@
 This metric gives us an aggregate of the complexity of documents, their sentences and words, measured by the average number of words per sentence and the average number of syllables per word.
-After obtaining an average for both hoax and real articles, we visualize these averages again in Figure 2, we find a median of 9.4 for real articles and 9.5 for hoax articles, again highlighting the similarities between these articles.
+After obtaining an average for both hoax and legitimate articles, we visualize these averages again in Figure 3, we find a median of 9.4 for legitimate articles and 9.5 for hoax articles, again highlighting the similarities between these articles.
+Timeline Analysis:
+Analysing the revision timelines of hoaxes and legitimate Wikipedia articles can give valuable insights into revision density and pattern that might differentiate the both sets of articles.
+To that end, we collect the timelines of hoax and legitimate articles (in all three hoax to legitimate ratio mentioned above) and add a timeline dataset to Hoaxpedia.
+Since some of the hoax articles were deleted from Wikipedia, so we are only able to obtain 164 hoax articles out of 311 in our dataset.
+With these hoax articles and their semantically similar legitimate counterparts, we construct the timeline dataset of Hoaxpedia.
+We check the timeline patterns through the use of Bayesian Online Changepoint Detection (BOCPD) Adams and MacKay (2007) with Kernel Density Estimation (KDE) Węglarczyk (2018) to obtain the dense regions for revision histories of hoax and legitimate articles.
+Figure 4 show a comparative timelines with marked normalized dense region for a hoax and a legitimate article as example.
+Timelines of hoax article show that they are quite dormant in terms of their number of revisions and their dense regions are mostly around the beginning and end of the article’s timeline, which can be attributed to the new page patrol and deletion discussion of hoaxes respectively.
+To test this hypothesis further, we divide the revision timelines of hoax and legitimate articles into quartiles (divide the timeline into four equal parts) and count the number of dense regions over total dense regions.
+The result for each quartile is given in Table 1, which shows that the proportion of dense region happening at the beginning and at the end are much higher for hoax articles than legitimate ones.
+We also show the normalised histogram distribution of hoax and legitimate (real) articles in Figure 5.
+(a) Plot for an example Hoax article
+(b) Plot for an example legitimate article
+Figure 4: Timeline based dense region plots for hoax and legitimate articles (with start and end lines for each region and normalized densities marked in grey)
+Quartile Hoax Real
+Q1 0.69 0.75
+Q2 0.02 0.17
+Q3 0.04 0.22
+Q4 0.75 0.42
+Table 1: Average distribution of dense regions per quartile (timeline divided into four parts) for hoax and legitimate (real) articles
+Figure 5: Histogram of normalized distribution for number of revisions in dense regions for hoax and legitimate (real) article
+Since we find difference in timeline patterns for hoax and legitimate articles, we also test the possibility of classifying hoax article using timeline only.
+We train a Support Vector Machine (SVM) Vapnik (2013) for this purpose, with TF-IDF Sparck Jones (1972) vectorizer fit on revision history binned by months (all revision timestamps converted to bins of MM-YYYY format) to build the timeline vocabulary.
+The results for Label 1 (Hoax) in F1-score is in Table 2, which shows that the classifier is able to perform well with discerning hoax article from legitimate ones, with an effect of data imbalance clearly showing in model’s performance.
+This further proves that revision history between hoaxes and legitimate articles are different in its pattern, and can be an important feature in detection of hoaxes.
+However, we also argue that timeline alone may not be able solve the detection of hoaxes, as it is a statistical feature prone to outliers.
+Data Split Precision Recall F1
+1H2R 0.86 0.91 0.88
+1H10R 0.89 0.78 0.83
+1H100R 0.97 0.69 0.80
+Table 2: Results of SVM Timeline classifier for Label 1 (Hoax) for all data splits
 The above analysis suggests that hoax Wikipedia articles are indeed well disguised, and so in the next section we propose a suite of experiments for hoax detection based on language models, setting an initial set of baselines for this novel dataset.
@@
 We cast the problem of identifying hoax vs.
-legitimate articles as a binary classification problem, in whcih we evaluate a suite of LMs, specifically: BERT-family of models (BERT-base and large Devlin et al.
+legitimate articles as a binary classification problem, in whcih we evaluate a suite of LMs, specifically: BERT family of models (BERT base and large Devlin et al.
 (2019), RoBERTa-base and large Liu et al.
@@
 This latter setting is interesting from a lexicographic perspective because it helps us understand if the Wikipedia definitions show any pattern that a model could exploit.
-Moreover, from the practical point of view of building a classifier that could dynamically “patrol” Wikipedia and flag content automatically, a definition-only model would be more interpretable (with reduced amiguity and focusing on core meaning/properties of the entity) and could have less parameters (handling smaller vocabularies, and compressed knowledge), which would have practical retraining/deployment implications in cost and turnaround.
+Moreover, from the practical point of view of building a classifier that could dynamically “patrol” Wikipedia and flag content automatically, a definition-only model would be more interpretable (with reduced ambiguity and focusing on core meaning/properties of the entity) and could have less parameters (handling smaller vocabularies, and compressed knowledge), which would have practical retraining/deployment implications in cost and turnaround.
 6 Results
@@
 We run a data ablation test on the full-text split of the dataset to find the impact of definition sentences.
-To this end, we remove the first sentence of each article, and replicate the “full text” classification experiment, focusing on RoBERTa-Large, the most consistent model.
+To this end, we remove the first sentence of each article, and run the fulltext classification experiment, focusing on RoBERTa-Large, the most consistent model.
 The results of this ablation experiment are shown in Table 4, suggesting that F1 goes down about 2% for the positive class when the definition sentence is missing.
@@
 Model Setting Precision Recall F1
-RoBERTaL Ft 1:2 0.83 0.80 0.82
-RoBERTaL Ft 1:10 0.82 0.71 0.76
-RoBERTaL Ft 1:100 0.67 0.51 0.58
-Table 2: Performance of RoBERTa-Large on Binary Classifications Without Definition Sentences in Articles (with Fulltext Hoax:Real ratio in Settings column)
+RoBERTaL 1H2R 0.83 0.80 0.82
+RoBERTaL 1H10R 0.82 0.71 0.76
+RoBERTaL 1H100R 0.67 0.51 0.58
+Table 4: Performance of RoBERTa-Large on Binary Classifications Without Definition Sentences in Articles (with Fulltext hoax to real ratio in Settings column)
+7 Large Language Models and HoaxPedia
+We explore the capabilities of open source Large Language Models (LLM) in detecting hoax articles through our proposed dataset.
+We select Llama2-7B and 13B Touvron et al.
+(2023), Llama3-8B Dubey et al.
+(2024), and Mistral-7B models Jiang et al.
+(2023) for the experiments, and the prompts used are given in the Appendix C.
+We take three experimental settings: perplexity, instruction-tuning, and supervised fine-tuning.
+The settings and their results are described below.
+Perplexity Experiments: We consider perplexity as an indicator for LLMs to predict distribution of of Hoax and legitimate articles, with the hypothesis that LLMs will have difficulty for predicting contents of hoax articles, resulting in higher perplexity.
+We test the LLMs in both definition and fulltext settings.
+The average perplexity results for both settings are shown in Figure 6, revealing that there is a significant difference between perplexity of hoax and legitimate articles in both settings.
+This suggests that LLMs struggle to predict distribution of Hoax articles.
+(a) Average perplexity scores for fulltext
+(b) Average perplexity scores for definition
+Figure 6: Average Model based perplexity scores for Fulltext and Definition Settings
+Prompting: For prompting, we consider zero-shot and few-shot prompt settings.
+The prompts are given in Appendix C, and the input setting are for both definition and fulltext.
+We report the results for Macro-F1 scores.
+The results are reported in Table 5, showing that the Mistral model show consistency in performance across the splits, albeit it is not the best model.
+Llama2-13B models perform the best for both the settings (definition and fulltext).
+The results also show that the performance difference between the definition and fulltext setting is marginal, as opposed to fine-tuned Language Models in Table 3.
+Model Name Zero-shot Few-shot
+Definition Fulltext Definition Fulltext
+Llama2-7B 0.48 0.50 0.51 0.52
+Llama2-13B 0.57 0.58 0.59 0.59
+Llama3-8B 0.33 0.40 0.35 0.40
+Mistral-7B 0.53 0.56 0.54 0.58
+Table 5: F1 score for Prompting experiment in Zero and Few shot setting for Definition and Fulltext splits
+Fine-tuning: We also fine-tune the models tested with perplexity and instruction tuning with HoaxPedia with supervised fine-tuning Touvron et al.
+(2023).
+The results of fine-tuning as Macro-F1 scores for both definition and fulltext setting are shown in Table 6.
+The results show significant improvement across all the settings for all the models.
+Llama3 shows most consistency across the settings, and also is the best model across the scenarios, with a performance improvement of more than 25% across the scenarios.
+However, the performance of the models in 1:100 hoax to legitimate scenario for all the models are around 50% F1 in both settings.
+Further analysis reveals that the models only predict the majority label, which can further attributed to imbalance in the dataset Powers (2020).
+A suitable weight adjusted training paradigm is required to handle the imbalance, but since it is not the scope of this work, we keep it for future work.
+Model Definition Fulltext
+1H2R 1H10R 1H100R 1H2R 1H10R 1H100R
+Llama2-7B 0.76 0.47 0.49 0.66 0.48 0.47
+Llama2-13B 0.80 0.48 0.50 0.60 0.63 0.50
+Llama3-8B 0.80 0.48 0.50 0.83 0.67 0.50
+Mistral-7B 0.71 0.55 0.49 0.68 0.53 0.49
+Table 6: Macro-F1 score for LLM Fine-tuning in degrees of data imbalance for Definition and Fulltext Settings
 8 Conclusion and Future Work
-We have introduced HoaxPedia, a dataset containing hoax articles extracted from Wikipedia, from a number of sources, from official lists of hoaxes, existing datasets and Web Archive.
-We paired these hoax articles with similar real articles, and after analyzing their main properties (concluding they are written with very similar style and content), we report the results of a number of binary classification experiments, where we explore the impact of (1) positive to negative ratio; and (2) going from the whole article to only the definition.
-This is different from previous works in that we have exclusively looked at the content of these hoax articles, rather than metadata such as traffic or longevity.
+We have introduced HoaxPedia, a dataset containing hoax articles extracted from Wikipedia, from a number of sources, from official lists of hoaxes, existing datasets, and the Web Archive.
+We paired these hoax articles with similar legitimate articles, and after analyzing their main properties (concluding they are written with very similar style and content), we report the results of a number of binary classification experiments, where we explore the impact of (1) positive to negative ratio; and (2) going from the whole article to only the definition.
+This is different from previous work in that we have exclusively looked at the content of these hoax articles, rather than metadata such as traffic or longevity.
 For the future, we would like to further refine what are the criteria used by Wikipedia editors to detect hoax articles, and turn those insights into a ML model, and explore other types of non-obvious online vandalism.
+9 Limitations
+We present a new dataset named Hoaxpedia and associated baselines from a wide variety of language models / large language models in this work.
+Our study shows that these types of dataset can be helpful in the area of free text disinformation detection.
+However, there are some limitations to our work that we aim to address here.
+The sets proposed here are small, with only 311 positive examples (hoaxes), which can be attributed to the fact that we only collect the examples that are explicitly labeled as hoaxes, rather than articles under discussion for hoaxes.
+Additionally, in our experiments, we do not conduct further investigation for model behaviors such as performance improvement of Longformer models in the hardest setting (cf.
+Section 6, implementing a weight-adjusted training paradigm for fine-tuning of LLMs (cf.
+Section 7) etc.
+We leave these analysis in future work, as the scope of this work is to introduce this dataset and establish the baseline results.
+Finally, we do not compare the results with existing work, mainly with Kumar et al.
+(2016), since the approaches mentioned in existing work are metadata dependent, and our approach is based on article text, we argue that the results may not be comparable.
 10 Ethics Statement
@@
 However, we acknowledge that deceitful content can be also used with malicious intents, and we will make it clear in any associated documentation that any dataset or model released as a result of this paper should be used for ensuring a more transparent and trustworthy Internet.
-Appendix A Appendix: Dataset Details
+Appendix A Dataset Details
 We release our dataset in 3 settings as mentioned in Section 5. The settings with data splits and their corresponding sizes are mentioned in Table 7.
@@
 Dataset Setting Dataset Type Split Non-hoax Hoax Total
-1Hoax2Real Definition Train 426 206 632
+1Hoax2legitimate Definition Train 426 206 632
 Test 179 93 272
-1Hoax2Real Full Text Train 456 232 688
+1Hoax2legitimate Full Text Train 456 232 688
 Test 200 96 296
-1Hoax10Real Definition Train 2225 203 2428
-Test 940 104 1044
-1Hoax10Real Full Text Train 2306 218 2524
-Test 973 110 1083
-1Hoax100Real Definition Train 20419 217 20636
-Test 8761 82 8843
-1Hoax100Real Full Text Train 22274 222 22496
-Test 9534 106 9640
+1Hoax10legitimate Definition Train 2,225 203 2,428
+Test 940 104 1,044
+1Hoax10legitimate Full Text Train 2,306 218 2,524
+Test 973 110 1,083
+1Hoax100legitimate Definition Train 20,419 217 20,636
+Test 8,761 82 8,843
+1Hoax100legitimate Full Text Train 22,274 222 22,496
+Test 9,534 106 9,640
 Table 7: Dataset details in different settings and splits
-Appendix B Appendix: Training Details
-We train our models with the configuration given below. We use one NVIDIA RTX4090, one NVIDIA V100 and one NVIDIA A100 GPU to train the models.
+Appendix B Language Model Training Details
+We train our Language Models with the configuration given below. We use one NVIDIA RTX4090 to train the LMs, one NVIDIA V100 and one NVIDIA A100 GPU to train the LLMs.
   • Learning rate: 2e-06
@@
   • Warm-up steps: 100
+Appendix C Prompt for LLM in-context learning
+The instruction prompt used for LLMs in their in-context learning with examples for few shot experiment are given below.
+You are a helpful knowledge management expert and excel at identifying whether an input Wikipedia article is a hoax or not.
+Wikipedia defines a hoax as ‘a deliberately fabricated falsehood made to masquerade as truth’.
+You take an Wikipedia article as input and return with the label citing hoax(Label 1) or real(Label 0) based only on the text of the article.
+Given an article from Wikipedia, your task is to analyze the article text to identify if the article is hoax or real.
+The Hoax and real articles are defined as follows: Hoax: An article that is deliberately fabricated falsehood made to masquerade as truth.
+Real: An article which contains information about an existing entity and are not fabricated.
+Your output should be a JSON dictionary with label that you found.
+Here are the possible labels with what they mean: 0 : The article is real article.
+1 : The article is a hoax article.
+Your input will be in the following format: INPUT: { Text: <Article text> } OUTPUT: { Label: <One of the label from the possible labels: 0 and 1, where 0 is real article and 1 is hoax article.> } Please respond with only the JSON dictionary containing label.
+You are instructed strictly to return output only in the format given above, nothing else.
+No yapping.
+Here are the examples used in few-shot experiments.
+Example 1: INPUT: { Text: Albion Dauti (born May 31, 1995 in Caracas, Venezuela) is a Venezuelan telenovela actor and presenter.
+Albion Dauti was born in Caracas, Venezuela.
+He studied acting at the Faculty of Arts in Caracas.
+In 2010 he started working as a presenter in Venevision.
+} OUTPUT: { Label: 1 } Example 2: INPUT: { Text: Michelle Madhok (born May 26, 1971) is the Founder and CEO of White Cat Media Inc.
+- DBA SheFinds Media, parent company of online shopping publication SheFinds.
+com and MomFinds.
+com.
+She writes a weekly style column for New York’s Metro newspaper and appears regularly on Fox News Channel, The Today Show, and The Tyra Banks Show.
+} OUTPUT: { Label: 0 }